OpenStudy (anonymous):

Can someone just help me with the steps... Passing through (3, 3) and perpendicular to y=1/3x + 7

OpenStudy (anonymous):

perpendicular tells us m1 x m2 = -1 m1=gradient of the known equation given, we can see it's 1/3 m2=gradient of the unknown equation which we want to find. 1/3 x m2 = -1 m2 = -3 now use the gradient found m2=3 and the point (3,3) to substitute into the formula to find an unknown equation. y-y1=m(x-x1) y-3=-3(x-3) y=............ (you can do this, I will check it for you)